Suppose the water at the top of Niagara Falls has a horizontal speed of 3.52 m/s just before it cascades over the edge of the falls. At what vertical distance below the edge does the velocity vector of the water point downward at a 78.3 ° angle below the horizontal?
Answer #1

Sol. :- We are given initial velocity along X-axis Vx= 3.2m/s

and let velocity along Y-axis = VY

So we have velocity V = Vx  + VY

So Vy should be such that it create angle with edge as shown in diagram

So Vy/Vx

So Vy = Vx

So Vy is around 17m/s .

We know that initial velocity along y axis is zero as water was moving only in forward direction just before falling from edge.

So using newton's equation of motion

V2 - U2 = 2aS , for vertical axis a = gravitational pull = 9.8m/s2 and U = 0

So S = V2/2g = = 14.74m

So at around 14.74 m of vertical height water will be at below the horizontal.

Upvote if helps.....